Tickets for a school concert were priced at $3 for students and $6 for non-students. There were 1,215 tickets sold for a total of $5,418.
How many student tickets were sold?
How many non-student tickets were sold?

Answer by jorel1380(3719)

m+n=1215
3m+6n=5418
3m+3n=3645
3n=1773
n=591
m=624
There were 624 student tickets sold and 591 non-student tickets..